### 26 Feb 2026 v1

# The DMA Streaming Framework: Kernel-Level Buffer Orchestration for High-Performance AI Data Paths

Abstract.
AI transport libraries move bytes efficiently, but they commonly assume that buffers are already correctly allocated, placed, shared, registered, and safe under completion and teardown pressure. This paper presents dmaplane, a Linux kernel module that makes this missing layer explicit as buffer orchestration. dmaplane exposes a stable kernel UAPI via /dev/dmaplane and composes ring-based command channels, DMA buffer lifecycle management, dma-buf export for cross-device sharing, a kernel-space RDMA engine, NUMA-aware allocation and verification, credit-based flow control, low-overhead observability, and GPU memory integration via PCIe BAR pinning. We evaluate orchestration sensitivity with measurements of NUMA cross-node penalties at DRAM scale, completion-safe flow control under sustained RDMA load, and GPU BAR mapping tiers versus cudaMemcpy. We also demonstrate end-to-end disaggregated inference by transferring KV-cache chunks between two machines using RDMA WRITE WITH IMMEDIATE and reconstructing tensor views on the receiver. RDMA measurements use Soft-RoCE; we distinguish measured results from provider-independent properties by construction.
conference: ; ; journalyear: 2026
1. Introduction
Modern AI systems are limited by host side data movement as often as by accelerator compute. Large language model workloads increasingly rely on point to point transfers between GPUs across hosts, including disaggregated inference where a prefill GPU produces KV cache state and a decode GPU consumes it. Transport systems can move bytes efficiently, but they commonly assume that buffers are already allocated on the correct NUMA node, shareable across devices without copies, registered for RDMA, and safe under completion pressure and teardown. These preconditions are not transport. They are buffer orchestration.
This paper presents dmaplane, a Linux kernel module that makes buffer orchestration an explicit kernel layer. dmaplane exposes a stable UAPI at /dev/dmaplane using ioctl and mmap plus optional dma-buf file descriptors, and composes ring based command channels, DMA buffer lifecycle management, dma-buf export for cross device sharing, a kernel space RDMA engine, NUMA aware allocation and verification, credit based flow control for completion safety, low overhead observability, and GPU memory integration via PCIe BAR pinning.
The repo source code is at https://github.com/marcoeg/dmaplane.
1.1. Goals and non goals
Goals. dmaplane has two goals:
  • Provide a coherent kernel subsystem for buffer lifecycle, placement, sharing, and flow control, integrated with dma-buf, NUMA placement, RDMA memory registration, and GPU BAR pinning constraints.
  • Expose a stable kernel UAPI, based on ioctl and mmap plus optional dma-buf file descriptors, that higher level transports can consume.
Non goals. dmaplane is not intended to provide:
  • Collective communication primitives or an NCCL replacement.
  • A serving framework, including batching, routing, prefix caching, or scheduler integration.
  • Connection management (rdma_cm), multi-tenant isolation, or production fault recovery.
1.2. Motivating workloads beyond disaggregated inference
Disaggregated inference is a primary motivation because it requires point to point KV cache movement with receiver notification and bounded completion pressure. The same buffer orchestration layer is required in several adjacent workloads.
Mixture of Experts inference. MoE dispatch and combine routes token batches to expert networks on different devices. Activations become fine grained point to point messages. Steady state throughput depends on staging buffer placement, repeated registration cost, and completion safety under bursty traffic.
Disaggregated training. Training pipelines increasingly separate data parallel and model parallel stages from optimizer and checkpoint services. Gradients, optimizer shards, and activation checkpoints move between devices and hosts without a single global collective barrier. Buffer lifecycle, cross device sharing, and teardown safety are required when multiple DMA engines touch the same buffers.
RL rollout weight transfer. Reinforcement learning systems commonly separate actors producing rollouts from learners updating weights. Periodic weight pushes from learners to many actors stress point to multipoint distribution. Orchestration determines whether weights can be staged once, shared, and transmitted efficiently without CPU copies and without completion overflow during fanout bursts.
LLM training with weight streaming. In weight streaming architectures, such as Cerebras weight streaming, host side pipelines stage large weight shards and stream them into the accelerator each step. The limiting factors are sustained bandwidth, locality to the correct NUMA node and PCIe root complex, and backpressure correctness. Buffer orchestration governs placement, sharing, and flow control under sustained streaming.
1.3. Contributions
This paper makes four contributions:
  1. (1)
    A concrete design and implementation of a kernel level buffer orchestration layer with explicit invariants for concurrency, lifecycle, and teardown (Sections 3 and 4).
  1. (1)
    Measured sensitivity results showing where orchestration choices dominate throughput and safety: NUMA placement effects at DRAM scale (Table 14), completion safe flow control under sustained load (Table 13), and GPU BAR mapping tier behavior versus cudaMemcpy (Table 15).
  1. (1)
    An end to end disaggregated inference demonstration that transfers KV cache chunks between two machines using RDMA WRITE WITH IMMEDIATE and reconstructs tensor views on the receiver (Section 5, Table 16).
  1. (1)
    A practical decomposition into independently testable subsystems with reproducibility guidance and appendices for build, deployment, and test mapping (Appendices A through C).
1.4. Paper organization
Section 2 summarizes required background and positions dmaplane relative to adjacent systems. Section 3 describes architecture and design invariants. Section 4 presents the core mechanisms in dependency order. Section 5 describes the disaggregated inference demo and protocol. Section 6 evaluates the system and discusses limitations. Section 7 concludes. References follow the main body.
2. Background and Related Work
dmaplane relies on existing Linux kernel primitives rather than reimplementing memory management or RDMA transport. This section states only the background required to follow the design and then positions dmaplane relative to neighboring systems.
2.1. Kernel primitives used by dmaplane
DMA allocation and mapping. Small control structures often use dma_alloc_coherent to obtain coherent memory and a DMA address for a specific device. Large data buffers typically use page backed allocation (alloc_pages and alloc_pages_node) and device specific mapping through scatter gather tables and the DMA mapping API.
NUMA placement. alloc_pages_node(node, ...) requests allocation from a specific NUMA node, but can fall back silently to other nodes. Correct placement therefore requires explicit verification after allocation.
dma-buf. The dma-buf framework enables zero copy sharing of DMA capable memory between kernel drivers via an exporter importer contract. A key constraint is that scatter gather tables must be constructed per importer attachment because DMA addresses depend on the importing device and its IOMMU context.
Kernel verbs RDMA. The InfiniBand verbs API in ib_core provides a provider independent interface for RDMA. Kernel consumers allocate PDs, CQs, QPs, and register memory regions. Kernel space ownership is required when the memory being registered is not userspace memory.
GPU VRAM integration requires a bridge across the struct page boundary and uses NVIDIA peer to peer pinning APIs, with teardown constraints enforced by an asynchronous unpin callback.
2.2. Positioning and closest neighbors
Table 1 positions dmaplane relative to systems that occupy adjacent layers.
Table 1. Capability positioning across the buffer orchestration and transport stack.
Capability dmaplane nvidia-peermem NCCL Mooncake TransferEngine
Buffer allocation (coherent, page-backed) Yes No internal No No
NUMA-aware placement Yes No topology detection mentioned mentioned
dma-buf export or import Yes importer only No No No
MR registration (kernel-space) Yes Yes (GPU BAR only) internal via libibverbs via libibverbs
Credit-based flow control Yes No per-channel credits implicit in batching per-connection
GPU BAR pinning and mapping Yes Yes via CUDA via CUDA via CUDA
Cross-machine RDMA transport loopback and peer QP No multi-transport multi-transport multi-transport
Kernel tracepoint instrumentation Yes No limited No No
Collective operations (AllReduce) No No Yes No No
Serving integration (batching, routing) No No No Yes No
NCCL. NCCL (NVIDIA Corporation, 2026a) implements GPU optimized collective communication and includes internal buffer pool and topology logic. It targets barrier synchronized collectives. dmaplane targets point to point pipelines and provides the kernel level buffer lifecycle and completion safety substrate beneath collective scheduling.
Mooncake and TransferEngine. Mooncake (Qin et al., 2024) and TransferEngine (Licker et al., 2025) provide transport abstractions across memory tiers and network providers in userspace. They assume that pinned memory, registration, and placement preconditions are satisfied. dmaplane externalizes and measures those preconditions by implementing kernel side allocation, sharing, and registration steps.
nvidia-peermem. nvidia-peermem (NVIDIA Corporation, 2026b) bridges GPU BAR pages into the RDMA stack by registering them as memory regions. It addresses a narrow GPU to NIC sharing path. dmaplane incorporates an equivalent pinning and teardown discipline as one subsystem and extends it with NUMA placement, dma-buf export, flow control, and observability.
3. Architecture and Design Invariants
dmaplane is a loadable Linux kernel module that registers one character device and exposes a stable UAPI via /dev/dmaplane. The device is a kernel service, not a PCIe device driver. It provides buffer orchestration primitives that other components can consume.
3.1. Architecture overview
Figure 1 shows the system structure. Userspace submits work through ioctl and maps buffers via mmap. Work executes inside per channel worker threads draining submission rings and posting completions to completion rings. Buffers are named objects referenced by IDs, allowing subsystems to compose without exposing raw pointers across the UAPI.
For receiver notification with remote placement, dmaplane uses RDMA WRITE WITH IMMEDIATE as illustrated in Section 5. Each operation produces a send completion on the sender and a receive completion on the receiver carrying a 32 bit immediate value. Each operation consumes one pre posted receive WR on the receiver.
Figure 1. dmaplane block diagram.
dmaplane block diagram.
The complete ioctl map, benchmark commands, and test mapping are provided in the appendices.
3.2. Locking model
dmaplane is concurrent. Multiple userspace threads can issue ioctls concurrently, and each channel has an independent worker thread. The module uses a strict lock ordering:
  • dev_mutex serializes global device operations that can sleep.
  • rdma_sem is a reader writer semaphore protecting the RDMA context, taking write mode for setup and teardown and read mode for fast paths.
  • buf_lock protects the buffer ID map and fast per buffer state transitions.
  • mr_lock protects the MR registry.
The ordering is dev_mutex, then rdma_sem, then buf_lock, then mr_lock.
3.3. Design invariants
The subsystems are coupled by invariants for lifetime, teardown, and safety.
  • Lock ordering invariant. Locks are acquired in one global order: dev_mutex, then rdma_sem, then buf_lock, then mr_lock. Enforcement: code paths follow the order and do not invert it. Failure prevented: deadlock under concurrent ioctls.
  • Teardown ordering invariant. debugfs entries are removed before device teardown, and RDMA teardown excludes in flight operations by taking rdma_sem in write mode. Enforcement: module exit order and semaphore usage. Failure prevented: use after free from debugfs reads or completion processing racing teardown.
  • mmap lifetime invariant. A buffer cannot be destroyed while it has active userspace mappings. Enforcement: mmap_count accounting on initial mmap plus VMA open and close callbacks. Failure prevented: freeing pages still mapped in a process VMA.
  • dma-buf exporter invariant. Scatter gather tables are constructed per importer attachment. Enforcement: dma-buf exporter map_dma_buf builds an SG table for the importing device. Failure prevented: invalid DMA addresses from wrong IOMMU context.
  • GPU unpin callback invariant. The GPU unpin callback must not block and must not run teardown under foreign locks. Enforcement: atomic revocation flag and deferred cleanup. Failure prevented: deadlock with NVIDIA driver internal locks.
  • Completion processing invariant. Completion polling is performed in controlled contexts and quiesced during teardown. Enforcement: polling design and teardown ordering. Failure prevented: asynchronous callbacks executing after resources are freed.
4. Core Mechanisms
This section presents the mechanisms that implement the buffer orchestration contract: command channels, buffer lifecycle and sharing, kernel RDMA ownership, completion safe flow control, and GPU integration constraints.
4.1. Channels and ring based dispatch
Each channel provides a submission ring and completion ring plus a worker thread. Userspace submits an entry, the worker executes the operation, and a completion entry returns status and metadata. Rings are fixed size circular buffers with head and tail indices protected by per ring spinlocks. Worker threads sleep on wait queues and wake on submission, and they stop via kthread_stop during teardown.
This mechanism provides a stable execution substrate for later subsystems, where the dominant costs come from DMA mapping, RDMA posting, or GPU BAR access rather than ring dispatch.
4.2. Buffer lifecycle, mmap, and dma-buf export
dmaplane supports two allocation paths:
  • Coherent allocation via dma_alloc_coherent for small control structures.
  • Page backed allocation via alloc_pages or alloc_pages_node for large data buffers that require NUMA steering, scatter gather, and MR registration.
Buffers are exposed to userspace without copies by mapping the backing pages into a VMA via vm_insert_page in a loop. Buffer teardown must prevent freeing memory while it remains mapped. dmaplane tracks mmap_count and rejects destroy if the count is non zero. A kernel detail matters: the VMA open callback does not run on the initial mmap call, so the initial mapping increments the count explicitly.
For cross device sharing, dmaplane exports buffers via dma-buf. The exporter provides callbacks to attach and map the buffer for each importing device. A key constraint is per importer scatter gather construction because DMA addresses depend on the importing device and its IOMMU context.
Figure 2. dma-buf export and per importer attachment mapping.
dma-buf export and per importer attachment mapping.
4.3. Kernel RDMA engine ownership
dmaplane owns RDMA resources in kernel space because it owns the pages being registered. Userspace verbs APIs register userspace pages, not kernel allocated pages.
The verbs resource chain is PD, CQs, QPs, and MRs. dmaplane allocates these in order and destroys them in reverse order, flushing QPs by transitioning to error state before destroy. IB_POLL_DIRECT is used for CQs so completions are polled explicitly rather than delivered via asynchronous callbacks.
Figure 3. RDMA resource hierarchy.
RDMA resource hierarchy.
4.4. Completion safe flow control
Completion queue overflow can discard completions and corrupt sender accounting. dmaplane enforces a credit invariant that bounds in flight operations by CQ capacity. Credits decrement on post and increment on completion poll, ensuring in_flight <= max_credits <= CQ_depth in the configured benchmark paths.
RDMA WRITE WITH IMMEDIATE also consumes receiver side receive WR capacity. A receiver pre posted receive window therefore acts as a second credit type. Safe operation requires bounding in flight WRITE WITH IMMEDIATE operations by both sender completion capacity and receiver notification capacity.
Figure 4. Timeline of send CQ credits and receive window credits in a WRITE WITH IMMEDIATE pipeline.
Timeline diagram showing a sender posting WRITE WITH IMMEDIATE chunks bounded by send credits and receive window credits, the receiver polling the receive completion queue and reposting receives, and the sender polling the send completion queue to reclaim credits.
4.5. GPU memory integration constraints
GPU VRAM does not provide struct page, so standard kernel DMA and RDMA mapping paths do not apply directly. dmaplane uses NVIDIA peer to peer pinning APIs to obtain BAR page descriptors and enforces a non blocking unpin callback contract by deferring cleanup outside the callback.
Host access to BAR mappings is sensitive to mapping type. Write combining mappings improve host to GPU write throughput relative to uncacheable mappings, while GPU to host reads remain limited by non posted PCIe read semantics. The measured tier behavior is reported in Table 15.
5. Disaggregated Inference Demonstration
This section demonstrates composition of the buffer orchestration layer by transferring KV cache state between two machines and running a decode phase on the receiver.
5.1. Pipeline overview
The demo separates inference into two roles:
  • Prefill machine. Runs tokenization and the forward pass to produce KV cache. Consolidates KV cache into a pinned GPU staging buffer and transfers it in fixed size chunks to the receiver.
  • Decode machine. Pre posts receives, accepts RDMA WRITE WITH IMMEDIATE notifications, reconstructs KV cache tensor views over the landing zone, and runs token generation.
The demo uses RDMA WRITE WITH IMMEDIATE so that each chunk is placed into a specific remote address range and also delivers a 32 bit immediate value identifying the chunk.
5.2. Protocol and flow control
Each KV cache transfer is chunked. For each chunk the sender posts IB_WR_RDMA_WRITE_WITH_IMM with:
  • local SGE referencing the staged chunk
  • remote address within the receiver landing zone MR
  • receiver rkey
  • immediate value encoding of layer_index and chunk_index (16-bit fields) as implemented in the accompanying artifact
On the receiver, each WRITE WITH IMMEDIATE consumes one pre posted receive WR and produces a receive completion carrying the immediate value. The receive window therefore provides receiver notification credits as defined in Section 4.4, and the sender side flow control provides send CQ credits. The combined bound applies because the verb produces completions on both sides.
A sentinel immediate value signals completion of the transfer, and the receiver verifies that expected chunks arrived before reconstruction.
5.3. Measured timing breakdown
Table 16 reports the measured pipeline timing breakdown for the two instance EC2 demo using Soft-RoCE.
Table 2. Disaggregated inference pipeline timing (two EC2 ‘g5.xlarge‘, Soft-RoCE).
Phase Time
Tokenization 1.2 ms
Prefill forward pass 45.3 ms
KV-cache consolidation 0.8 ms
KV-cache transfer 52.1 ms
KV-cache reconstruction 0.003 ms
Time-to-first-token (TTFT) 98.2 ms
Decode throughput 45.3 tok/s
Decode latency (per token) 22 ms average
6. Evaluation and Discussion
This section presents four result sets that support the buffer orchestration thesis: flow control safety under sustained RDMA load (Table 13), NUMA placement sensitivity at DRAM scale (Table 14), GPU BAR tier behavior versus cudaMemcpy (Table 15), and the end to end demo timing breakdown (Table 16). All other microbenchmarks, extended tables, and setup details are provided in the appendices.
6.1. Flow control under sustained streaming
Table 13 reports sustained streaming and stress results under Soft-RoCE loopback. The reported configurations include a high credit configuration and a low credit stress configuration that generates large numbers of stalls while reporting zero CQ overflows.
Table 3. RDMA streaming and flow control results (Soft-RoCE loopback).
Metric Value
Sustained throughput (10 s, max_credits=64) 1,037 MB/s average
Per-second window range 1,047 to 1,087 MB/s
Window spread 3.8%
CQ overflows 0
Credit stalls 0
Stress configuration (max_credits=4, high=3, low=1)
Credit stalls (5 s) 72.7 million
CQ overflows 0
These results are consistent with the invariant in_flight <= max_credits <= CQ_depth for the signaled work request configuration used in the evaluation.
6.2. NUMA placement sensitivity
Table 14 reports cross node CPU memcpy bandwidth on AWS c5.metal at 1 MB and 64 MB. The 1 MB case does not show a cross node penalty, while the 64 MB case does, consistent with cache hiding effects at small working sets.
Table 4. Cross node memcpy bandwidth on AWS ‘c5.metal‘. Values are MB/s.
Buffer size Node 0 to 0 Node 0 to 1 Node 1 to 0 Node 1 to 1 Cross-node penalty
1 MB 13,284 13,218 12,976 13,051 < 1% (fits in cache)
64 MB 6,778 5,577 5,013 6,095 18% symmetric
Methodology: based on the accompanying artifact; iteration count and warmup policy are not specified in the paper.
The practical implication for buffer orchestration is that placement errors can be silent and can appear only at DRAM scale buffer sizes.
6.3. GPU memory access tiers
Table 15 consolidates measured tiers for GPU BAR access and cudaMemcpy on RTX 5000 Ada, plus measured software RDMA loopback and cross machine behavior from the accompanying artifact.
Table 5. GPU memory access hierarchy (RTX 5000 Ada). Cross-machine values are over 1 Gigabit Ethernet to a remote NUC.
Access method Host to GPU (MB/s) GPU to host (MB/s) Section
UC BAR (ioremap) 44 6 Accompanying artifact
WC BAR (ioremap_wc, 4 MB) 10,097 107 Accompanying artifact
cudaMemcpy (16 MB) 12,552 13,124 Accompanying artifact
GPU RDMA loopback (rxe, 64 KB) n/a ~20 Accompanying artifact
GPU RDMA cross-machine (1 Gigabit) n/a 11.6 Accompanying artifact
Derived: 64 KB / 3.2 ms = ~20 MB/s effective for the loopback entry, based on the
measured 64 KB loopback latency.
Methodology: BAR and cudaMemcpy tiers are mean of three runs per size; cross machine tiers are from the two machine configuration described in the paper.
The NIC DMA engine BAR read tiers described in NVIDIA documentation are not measured in this work and must be treated as external claims when referenced.
6.4. End to end demo
Table 16 provides the end to end timing breakdown for the disaggregated inference demo. The largest component in time to first token in the reported run is the KV cache transfer time under Soft-RoCE and Ethernet.
6.5. Limitations and generality
6.5.1. Measured on Soft-RoCE
RDMA measurements in this paper run over Soft-RoCE, a software provider that executes verbs in kernel software and uses CPU memcpy. Results reflect this provider behavior and host CPU scheduling effects.
6.5.2. Provider independent by construction
dmaplane uses provider independent kernel verbs APIs and standard completion semantics. The flow control invariant is expressed over completion accounting rules, and the dma-buf sharing model follows the kernel framework contract. These properties do not depend on a specific RDMA NIC provider.
6.5.3. Not validated on hardware RDMA
The paper does not include performance measurements on hardware RDMA NICs such as ConnectX or EFA. GPU pinning uses the same peer to peer API used in GPUDirect RDMA environments, but NIC BAR DMA tiers are not measured in this work.
Implementation details, scripts, and additional measurements are provided in the accompanying artifact and appendices.
7. Conclusion
This paper presents dmaplane, a Linux kernel module that implements buffer orchestration as a first class layer beneath transport and above devices. dmaplane provides a stable kernel UAPI and composes mechanisms for buffer lifecycle, cross device sharing, kernel RDMA ownership, completion safe flow control, NUMA aware placement, observability, and GPU BAR pinning constraints.
Four measured result sets support the framing:
  • NUMA placement penalties can be hidden at cache scale and appear at DRAM scale (Table 14).
  • Credit based flow control prevents CQ overflow under sustained load in the reported configurations (Table 13).
  • BAR mapping tier choice changes host to GPU throughput dramatically while GPU to host reads remain limited (Table 15).
  • The end to end demo composes these primitives to transfer KV cache and run decode on the receiver (Table 16).
Future work includes validation on hardware RDMA providers and integration paths where dmaplane is consumed by higher level transports or serving frameworks without changing the kernel layer contract.
Appendix A Methodology
This appendix provides (1) a phase-to-subsystem mapping retained as internal project terminology and (2) a development process note describing AI-assisted development.
A.1. Development process note
The phase structure was used to bound implementation scope, define interfaces, and isolate test surfaces. dmaplane was built using Claude Code as a collaborative tool for implementation, debugging, and test suite generation. The bounded scope of each phase, with a clear specification, a defined interface to previous phases, and an isolated test surface, supported iterative development and verification.
Appendix B Build and Deployment
This appendix consolidates build and deployment details referenced throughout the paper. The repo source code is at https://github.com/marcoeg/dmaplane
B.1. Kernel and configuration requirements
The work targets a stock Ubuntu kernel (6.5). Required kernel facilities include:
  • RDMA core (ib_core) and the kernel verbs interface
  • Soft-RoCE provider (rdma_rxe) for the software provider measurements
  • debugfs (CONFIG_DEBUG_FS) for debugfs inspection
  • ftrace and tracepoints for optional runtime tracing
If debugfs is not mounted or CONFIG_DEBUG_FS is disabled, debugfs initialization returns silently and is not treated as fatal.
B.2. Module build and load
The module is built out of tree using Kbuild, the Linux kernel build sytstem. When tracepoints are enabled, the build must satisfy TRACE_INCLUDE_PATH and include path requirements so define_trace.h can re-include the trace header during macro expansion.
Loading uses standard module tooling:
sudo insmod dmaplane.ko ls -l /dev/dmaplane
Unloading requires that userspace closes the device file and releases mappings so teardown invariants are satisfied:
sudo rmmod dmaplane
B.3. Soft-RoCE bring-up (rxe)
Soft-RoCE is used for the RDMA provider in the measurements. A typical bring-up sequence:
sudo modprobe rdma_rxe sudo rdma link add rxe_eth0 type rxe netdev <netdev> ls /sys/class/infiniband
The RDMA device name (for example, rxe_eth0) is passed to dmaplane through the RDMA setup ioctl.
B.4. debugfs mount and inspection
Mount debugfs if not already mounted:
sudo mount -t debugfs none /sys/kernel/debug
Inspect driver state:
cat /sys/kernel/debug/dmaplane/stats cat /sys/kernel/debug/dmaplane/buffers cat /sys/kernel/debug/dmaplane/rdma cat /sys/kernel/debug/dmaplane/flow cat /sys/kernel/debug/dmaplane/histogram
B.5. Tracepoint enablement
Enable tracepoints via ftrace:
echo 1 > /sys/kernel/debug/tracing/events/dmaplane/dmaplane_rdma_post/enable echo 1 > /sys/kernel/debug/tracing/events/dmaplane/dmaplane_rdma_completion/enable cat /sys/kernel/debug/tracing/trace_pipe
B.6. NVIDIA driver constraints and runtime symbol bridge
[…]
Appendix C Test and Observability Map
This appendix summarizes the test surface and observability mechanisms described in the paper and referenced by the main body. Where specific filenames are not provided in the accompanying artifact, entries are marked as not specified.
C.1. Test suite map
Table C.1. Test suite map by subsystem.
Subsystem Test entry point Pass condition
Device and ioctl dispatch not specified Module loads, /dev/dmaplane present, ioctls return expected codes.
Rings and workers userspace stress harness No kernel warnings, no lockdep issues, no data corruption, clean unload.
DMA buffers and mmap not specified Create buffer, mmap, write and verify, munmap, destroy succeeds.
dma-buf export not specified Export returns dma-buf fd, attach and detach paths do not leak, release callback fires.
RDMA setup and SEND and RECV ioctl benchmarks QP setup succeeds, SEND and RECV completes, CQs polled without errors.
MR registration ioctl MR tests MR registration returns lkey and rkey when requested, deregistration releases mappings.
NUMA topology IOCTL_QUERY_NUMA_TOPO Returns distance matrix and node inventory without kernel warnings.
NUMA memcpy benchmark IOCTL_NUMA_BENCH Produces per-cell bandwidth values without kernel warnings.
Flow control IOCTL_SUSTAINED_STREAM Zero CQ overflows under configurations in Table 3.
QD sweep IOCTL_QD_SWEEP Produces throughput and latency values (Table 4).
Instrumentation debugfs inspection debugfs files readable when enabled, histogram format produced during streaming.
GPU pin and BAR mapping IOCTL_GPU_PIN and IOCTL_GPU_BENCH Pin succeeds when NVIDIA driver present, unpin callback safe.
GPU RDMA loopback not specified Loopback transfer verifies bytes match, no kernel warnings.
WRITE WITH IMMEDIATE helpers IOCTL_RDMA_WRITE_IMM and related ioctls Receiver polls imm_data and receives expected identifiers.
Disaggregated inference prefill and decode scripts Demo produces coherent output and timings consistent with Table 2.
Methodology: derived from the paper text and artifact descriptions; specific script and binary names are not fully specified in the paper.
C.2. Observability summary
dmaplane provides two observability paths:
  • Tracepoints. Optional kernel tracepoints for events such as RDMA posting, completions, and flow stalls. When disabled, tracepoints compile to near no op behavior.
  • debugfs. Read-only debugfs files under /sys/kernel/debug/dmaplane/ support live inspection of counters, buffers, RDMA state, flow state, and the latency histogram. The histogram output format is described in this paper.
Appendix D Glossary
ACPI:
Advanced Configuration and Power Interface. Firmware interface that provides tables used by the OS, including NUMA topology.
BAR:
Base Address Register. PCIe mechanism that maps a device memory region into the host physical address space.
BAR1:
NVIDIA term for the PCIe BAR aperture that exposes portions of GPU VRAM to the host.
CAS:
Compare and swap. Atomic primitive used to update shared values without locks.
CBQ:
Completion queue. See CQ.
CQ:
Completion Queue. RDMA data structure that receives Work Completions for completed operations.
CQE:
Completion Queue Entry. A single completion record placed into a CQ.
CPU:
Central Processing Unit.
CUDA:
NVIDIA Compute Unified Device Architecture. Platform and programming model for NVIDIA GPUs.
DMA:
Direct Memory Access. Mechanism where a device reads or writes memory without CPU copying.
DRAM:
Dynamic Random Access Memory. System main memory.
EFA:
Elastic Fabric Adapter. AWS network interface supporting high performance communication, including RDMA like semantics.
ENA:
Elastic Network Adapter. AWS virtual network interface used on many EC2 instances.
GID:
Global Identifier. RDMA addressing identifier used in RoCE and InfiniBand.
GPU:
Graphics Processing Unit.
HCA:
Host Channel Adapter. RDMA capable network adapter used in InfiniBand and RoCE environments.
HBM:
High Bandwidth Memory. GPU attached memory (for example HBM2e) with very high bandwidth relative to DRAM.
IB:
InfiniBand. RDMA interconnect architecture and related kernel subsystem interfaces.
IOMMU:
Input Output Memory Management Unit. Hardware that translates device DMA addresses to physical memory and enforces access control.
KV:
cache Key value cache. Transformer attention state produced during prefill and consumed during decode.
LID:
Local Identifier. InfiniBand subnet local address.
LLC:
Last Level Cache. Typically the largest shared CPU cache (often L3).
LLM:
Large Language Model.
MR:
Memory Region. RDMA object authorizing NIC access to a range of memory, identified by lkey and rkey.
NIC:
Network Interface Card.
NUMA:
Non Uniform Memory Access. System architecture where memory is divided into nodes with different access costs depending on CPU locality.
P2P:
Peer to peer. In this paper, typically refers to GPU peer to peer memory pinning and access for third party devices.
PD:
Protection Domain. RDMA resource that scopes access permissions for QPs, CQs, and MRs.
PFC:
Priority Flow Control. Ethernet mechanism used to reduce packet loss in RoCE deployments.
QP:
Queue Pair. RDMA endpoint containing a Send Queue and Receive Queue.
RC:
Reliable Connected. RDMA transport type that provides reliable in order delivery.
RDMA:
Remote Direct Memory Access. Mechanism enabling one host to read or write memory on another host without remote CPU involvement in the data path.
RECV:
WR Receive Work Request. RDMA receive queue entry posted by the receiver to accept incoming SEND or WRITE WITH IMMEDIATE notifications.
SG:
Scatter gather. Representation of a buffer as a list of segments, often page sized, used for DMA mapping.
SGE:
Scatter Gather Element. One segment descriptor referenced by an RDMA work request.
SGT:
Scatter Gather Table. Collection of SG entries describing a buffer.
SLIT:
System Locality Information Table. ACPI table describing relative NUMA node distance.
SQ:
Send Queue. RDMA queue inside a QP that holds outgoing work requests.
SOSP:
ACM Symposium on Operating Systems Principles.
TLP:
Transaction Layer Packet. PCIe transaction unit for reads and writes.
TTFT:
Time to first token. Time from request start to the first generated token in an inference pipeline.
UPI:
Ultra Path Interconnect. Intel inter socket interconnect.
VMA:
Virtual Memory Area. Kernel structure representing a contiguous memory mapping in a process.
VRAM:
Video RAM. GPU local memory.
WC:
Write combining. CPU memory mapping type that buffers and coalesces writes, used for BAR mappings via ioremap_wc.
WR:
Work Request. RDMA operation descriptor posted to a QP, such as SEND or RDMA WRITE.